In a random sample of 25 students, 14 students answered "Yes" to the question “Do you use a grocery delivery app?” Use the “plus-four” method to find a 95% confidence interval for the true proportion of students who use a grocery delivery app. Round to 3 decimal places, if necessary.
Answer #1

Using plus four method we need to 2 in the number of students who answered yes and 4 in the sample size. So we have:

X = 14+2=16 , n = 25+4=29

Here we have x -16, n=29,p =0.5517 Level of significance: α-0.05 Critical value: 1.96 The standard deviation is: sqrt((0.5517